# Less Is More [Uncaptioned image]: Scalable Visual Navigation from Limited Data

### v1

@@ I Introduction
Scaling-first approaches have proven particularly effective in autonomous driving, where virtually unlimited human demonstration data can be collected. However, in many real-world robotic settings, data is scarce, necessitating the incorporation of more first-principled approaches into the training of navigation policies. In such settings, transparent and rigorous analysis of model limitations becomes essential for guiding future development. Under this lens, we identify five core limitations that can pose significant challenges for current visual navigation models:
(i) The navigation capability relies on costly real-world expert demonstrations. (ii) Action diversity in existing datasets is limited, biasing navigation toward simple, conservative motion. (iii) Models must learn geometric obstacle avoidance entirely from demonstrations. (iv) Current models trained on cross-embodiment data cannot choose actions that are desirable for a specific embodiment. (v) Training often focuses only on feasible goals. In response, we ask: How can we build upon the promising imitation learning framework while efficiently supervising navigation policies at scale to acquire both semantic understanding and robust navigation behavior, conditioned on the embodiment?
@@
Across held-out GrandTour missions, LiMo trained on the augmented dataset (teleop + geometric) achieves substantially higher Success weighted by (normalized inverse) Path Length(SPL) [3] than when trained solely on teleop-data, while preserving strong performance on teleop-style splits. We show that the policy adapts paths to visible structures and selects semantically reasonable shortcuts when safe, while transparently exposing the approach’s current limitations. We further demonstrate real-robot deployment, illustrating LiMo’s practicality as a drop-in replacement for local path planning. We release all datasets and pretrained models to enable reproducibility and further research.
@@ II-A Geometric Path Planning
Existing geometric planners either use raw sensory data directly or rely on aggregated spatiotemporal geometric information, such as elevation maps, to compute feasible and safe paths [8]. These planners typically frame path planning as a search or optimization problem. Prominent approaches include graph-based search algorithms [14] and optimization methods minimizing a cost function, such as MPPI [46] or iCEM [35].
@@ II-B Visual Navigation
Visual navigation aims to learn navigation behaviors directly from camera input. Previous works leverage reinforcement learning [50, 29] and imitation learning (IL) [12, 26] to train visuomotor control policies from raw images.
IL has produced remarkable advances in autonomous driving (AD)  [23, 6, 24, 30]. Lately, several large-scale, high-quality AD datasets have been made available [48, 43, 13]. The use of IL for navigating other types of mobile robots is hindered by the limited availability of demonstration data. Collecting expert demonstrations is costly, and existing datasets typically cover only a few hours of operation, which limits robustness and generalization.
Recent goal-image–conditioned models, such as ViNT [39], use transformer-based architectures to learn navigation policies intended to generalize across varied settings and robot embodiments. NoMaD [41] extends ViNT by incorporating a diffusion head for trajectory prediction, which supports goal-conditioned navigation and unconditioned exploration. FlowNav [11] is based on NoMaD but replaces diffusion with conditional flow matching (CFM) to improve inference efficiency. Additionally, FlowNav uses an off-the-shelf monocular depth estimation model to generate depth priors, which are then fed to the model. These models use a goal image (the observation when the robot is at the goal) for goal-conditioning. This goal image is available during training on pre-recorded robot demonstrations, but it may not be available during deployment.
To overcome the limited data problem, [39, 41, 11] train on a mix of datasets; however, the overall scale remains limited. The total hours reported for ViNT/NoMaD training are $80\text{\,}\mathrm{h}$ across the collected datasets (see Tab I) and even less for FlowNav, which is modest given the complexity of the task. These works compellingly demonstrate the promise of IL for navigation; nevertheless, the scarcity of expert data continues to hinder generalization.
@@
LiReN [42] approaches data scarcity by continuously improving the model during deployment. They first use offline RL to train a navigation policy on the same multi-embodiment dataset as ViNT. Then they deploy the model and use an actor-critic algorithm to continuously improve it during deployment without needing a human expert to collect more data. LeLaN [16] uses NoMaD and a vision–language model (VLM) to automatically label first-person view (FPV) YouTube videos for navigation policy learning.
CityWalker [25] also uses a dataset of FPV videos from YouTube, on which it runs visual odometry (VO) to automatically generate supervision. The FrodoBots-2k dataset [17] uses crowdsourcing; volunteers remotely drive inexpensive robots (e.g., Earth Rover Zero) to collect demonstrations in urban pedestrian zones. This strategy scales data collection without on-site experts; however, the dataset suffers from poor tracking, inaccurate localization, and low-quality sensors. This is addressed by using Model-Based ReAnnotation (MBRA) [17], an approach that transforms a low-quality dataset into a filtered version with high-quality supervision.
@@
OmniVLA [15] is an omni-modal vision-language-action (VLA) model for navigation that supports goal conditioning via 2D goal poses, egocentric goal images, natural language instructions, and their combinations. Built on top of OpenVLA [22], OmniVLA leverages a large pretrained VLA backbone and is further trained on approximately $9500\text{\,}\mathrm{h}$ of navigation data. This dataset is the combination of [39, 41, 16, 17, 47] datasets from 10 different robotic platforms. Notably, the dataset mixture is heavily skewed towards BDD-V, i.e., on-road autonomous driving data. They use a customized MBRA-style [17] approach to generate synthetic, feasible supervision for their target robots on BDD-V.
@@ III-B MPPI Planning
$\min_{\mathbf{a}_{1:N}}\quad J(\mathbf{s}_{1},\mathbf{a}_{1:N}),$ (1)
where $\mathbf{a}_{1:N}$ is a sequence of command velocities $\mathbf{a}_{i}=[v_{i}^{x},v_{i}^{y},\omega_{i}]$, with $v_{i}^{x},v_{i}^{y},\omega_{i}$ representing the linear and angular velocities, respectively, and $J:(\mathbf{s}_{1},\mathbf{a}_{1:N})\mapsto c\in\mathbb{R}$ denoting a predefined cost function that encourages goal-reaching and safety, conditioned on the current state $\mathbf{s}_{1}\in\mathrm{SE}(2)$.
MPPI solves the optimization by iteratively sampling a population of command velocity sequences $\mathbf{a}_{1:N}$. It then uses $J$ to weight the population members and update the population distribution, preferring low-cost members.
The future states are computed based on the current state $\mathbf{s}_{1}$ and a dynamics model $\mathbf{s}_{i+1}=f(\mathbf{s}_{i},\mathbf{a}_{i})$. The dynamics model $f$ can be learned or classically derived. In our case, we use a simple piecewise-constant velocity kinematic model in $\mathrm{SE}(2)$:
@@
$\begin{bmatrix}x_{i+1}\\ y_{i+1}\end{bmatrix}$ $=\begin{bmatrix}x_{i}\\ y_{i}\end{bmatrix}+\mathbf{R}(\theta_{i})\begin{bmatrix}v_{x}^{i}\\ v_{y}^{i}\end{bmatrix}\Delta t,$ (3)
@@
$\begin{split}J(\mathbf{s}_{1},\mathbf{a}_{1:N})&=w_{\text{trav}}C_{\text{trav}}(\mathbf{s}_{1:N})+w_{\text{goal}}C_{\text{goal}}(\mathbf{s}_{1:N})\\ &\quad+w_{\text{effort}}C_{\text{effort}}(\mathbf{a}_{1:N}).\end{split}$ (4)
@@ III-B1 Traversability Cost ($C_{\text{trav}}$)
A convolutional neural network (CNN) [28, 7] is used to compute a traversability map $\mathbf{T}$ from the elevation map $\mathbf{E}$. The CNN is trained to provide an embodiment-specific traversability estimate based on the robot’s capabilities. The traversability cost is the sum of the grid cells of $\mathbf{T}$ traversed by the robot along $\mathbf{s}_{1:N}$.
@@ III-B3 Effort cost ($C_{\text{effort}}$)
This is the weighted sum of the linear, lateral, and angular velocities $w_{\text{lin}}|v^{x}|+w_{\text{lat}}|v^{y}|+w_{\text{ang}}|\omega|$ over the command sequence, encouraging efficient motion while discouraging unnecessary sideways motion and excessive turning.
@@ III-C1 Teleoperation Dataset ($\mathcal{D}_{\text{TEL}}$)
$\mathcal{D}_{\text{TEL}}$ consists of fixed-horizon, goal-conditioned segments extracted from the recorded teleoperated trajectories and serves to transfer semantic preferences and motion characteristics to the policy. Implementation details are provided in Sec. -A1.
@@ III-C2 Geometric Dataset ($\mathcal{D}_{\text{GEO}}$)
To increase the diversity and scale of the training data, we employ the previously introduced MPPI-planner to generate additional demonstrations. For every camera frame $\mathbf{I}$, we independently sample $K$ random goals $\mathbf{g}_{1},...,\mathbf{g}_{K}$ from a Gaussian distribution in the robot-centric coordinate frame (see Tab III in the Appendix). Given the matching elevation map $\mathbf{E}$ and the goals, MPPI plans the paths $\mathbf{s}_{1:N}^{1},...,\mathbf{s}_{1:N}^{K}$; generating $K$ diverse training samples $(\mathbf{I},\mathbf{g}_{i},\mathbf{s}_{1:N}^{i})$ that are saved in $\mathcal{D}_{\text{GEO}}$.
The MPPI planner reliably generates geometrically feasible paths by leveraging privileged information from accurate elevation mapping. This enables the creation of demonstrations that are more diverse and exploratory than those obtained from real-world teleoperation alone. Even when a randomly sampled goal is infeasible, the planner still produces safe, embodiment-consistent behavior, either progressing toward the goal in a best-effort manner or exploring alternative routes while avoiding obstacles.
@@ III-D Policy Architecture
Figure 3: Overview of LiMo’s architecture. The policy takes as input a single RGB image $\mathbf{I}$, and a goal pose $\mathbf{g}=(x,y,\theta)$ in the robot-centric frame. Image features are extracted using a frozen DINOv2 encoder and combined with learned positional embeddings. A transformer decoder conditioned on the goal embedding predicts NN waypoint embeddings, which are linearly projected to $N$ waypoints $(x,y,\theta)$, forming the output trajectory.
@@ IV-A Datasets
The experiments build on the GrandTour dataset [10], which comprises $6\text{\,}\mathrm{h}$ of highly diverse deployments of an ANYmal D robot across 49 sites in Switzerland, with varying weather and lighting conditions. From these missions, we derive three datasets, in accordance with Secs. III-C1 and III-C2: (i) the teleoperation dataset $\mathcal{D}_{\text{TEL}}$, (ii) the geometric-planner-labeled dataset $\mathcal{D}_{\text{GEO}}$ with $K=10$ trajectories per image, each with $N=50$ waypoints evenly spaced over $T=$ $5\text{\,}\mathrm{s}$, and finally, (iii) the augmented dataset $\mathcal{D}_{\text{AUG}}=\mathcal{D}_{\text{TEL}}\cup\mathcal{D}_{\text{GEO}}$. For evaluation, we reserve six GrandTour missions exclusively as test scenes (see Tab III in the Appendix for further details on dataset size).
@@ IV-C1 Metrics
Following the recommendations of [3], we evaluate trajectories on held-out missions using geodesic distances (GD) and Success weighted by normalized inverse Path Length (SPL). A path is deemed successful if (i) the GD, derived from the traversability map, of the final waypoint to the goal is $\leq$1.0\text{\,}\mathrm{m}$\$ and (ii) the robot does not collide. Let $S_{i}\in\{0,1\}$ indicate success for sample $i$, $\ell_{i}$ be the geodesic start-to-goal distance, and $p_{i}$ the executed path length; then the SPL is defined as:
$\mathrm{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\,\frac{\ell_{i}}{\max(p_{i},\ell_{i})}\,.$ (5)
We evaluate only goals reachable by the geometric planner, to avoid penalizing methods for failing to reach inherently infeasible goals.
@@ IV-C2 Baselines
Straight line
Paths go straight from the robot to the goals, capped at $5\text{\,}\mathrm{m}$ in length.
Geometric planner
The path obtained by running the planner described in Sec. III-B.
@@ IV-C3 Results
OmniVLA (open-loop)[15] $12.6$ $67.2$ $41.5$
@@
Straight line paths $12.5$ $87.5$ $87.5$
Real world paths $11.0$ $89.0$ $87.5$
@@
OmniVLA (open-loop)[15] $14.2$ $46.5$ $33.5$
@@
Straight line paths $23.1$ $76.9$ $76.9$
@@
As shown in Tab II, incorporating geometric demonstrations yields substantial SPL improvements on the diverse $\mathcal{D}_{\text{AUG}}$ test set, indicating that geometric supervision significantly strengthens performance in complex, unseen scenarios (Q1, Q2). As expected, when evaluating only on goals drawn from the $\mathcal{D}_{\text{TEL}}$ distribution, the improvement is smaller.
@@
We observe that OmniVLA achieves competitive collision avoidance, even on the more challenging $\mathcal{D}_{\text{AUG}}$ test set, but exhibits reduced goal-reaching performance under our open-loop evaluation. This behavior is expected in part due to the evaluation protocol: OmniVLA is trained for closed-loop replanning with fixed-length waypoint chunks rather than open-loop trajectory termination at a specified goal.

### v2

@@ Abstract
Project page: https://leggedrobotics.github.io/less-is-more
@@ I Introduction
Scaling-first approaches have proven particularly effective in autonomous driving, where virtually unlimited human demonstration data can be collected. However, in many real-world robotic settings, data is scarce, necessitating the incorporation of more first-principled approaches into the training of navigation policies. In such settings, transparent and rigorous analysis of model limitations becomes essential for guiding future development.
Under this lens, we identify five core limitations that can pose significant challenges for current visual navigation models: (i) The navigation capability relies on costly real-world expert demonstrations. (ii) Action diversity in existing datasets is limited, biasing navigation toward simple, conservative motion. (iii) Models must learn geometric obstacle avoidance entirely from demonstrations. (iv) Current models trained on cross-embodiment data cannot choose actions that are desirable for a specific embodiment. (v) Training often focuses only on feasible goals.
In response, we ask: How can we build upon the promising imitation learning framework while efficiently supervising navigation policies at scale to acquire both semantic understanding and robust navigation behavior, conditioned on the embodiment?
@@
Across held-out GrandTour missions, LiMo trained on the augmented dataset (teleop + geometric) achieves substantially higher Success weighted by normalized inverse Path Length (SPL) [3] than when trained solely on teleop-data, while preserving strong performance on teleop-style splits. We show that the policy adapts paths to visible structures and selects semantically reasonable shortcuts when safe, while transparently exposing the approach’s current limitations. We further demonstrate real-robot deployment, illustrating LiMo’s practicality as a drop-in replacement for local path planning. We release all datasets and pretrained models to enable reproducibility and further research.
@@ II-A Geometric Path Planning
Existing geometric planners either use raw sensory data directly or rely on aggregated spatiotemporal geometric information, such as elevation maps, to compute feasible and safe paths. These planners typically frame path planning as a search or optimization problem. Prominent approaches include graph-based search algorithms [11] and optimization methods minimizing a cost function, such as MPPI [41] or iCEM [30].
@@ II-B Visual Navigation
Visual navigation aims to learn navigation behaviors directly from camera input. Previous works leverage reinforcement learning (RL) [45] and imitation learning (IL) [23] to train visuomotor control policies.
IL has produced remarkable advances in autonomous driving (AD)  [20, 21, 26]. Lately, several large-scale, high-quality AD datasets have been made available [43, 38]. In contrast, the use of IL for navigating other types of mobile robots is hindered by the limited availability of demonstration data. Collecting expert demonstrations is costly, and existing datasets typically cover only a few hours of operation, which limits robustness and generalization.
Recent goal-image–conditioned models, such as ViNT [34], use transformer-based architectures to learn navigation policies intended to generalize across varied settings and robot embodiments. NoMaD [36] extends ViNT by incorporating a diffusion head for trajectory prediction, which supports goal-conditioned navigation and unconditioned exploration. FlowNav [10] is based on NoMaD but replaces diffusion with conditional flow matching to improve inference efficiency. Additionally, FlowNav uses an off-the-shelf monocular depth estimation model to generate depth priors, which are then fed to the model. These models use a goal image (the observation when the robot is at the goal) for goal-conditioning. This goal image is available during training on pre-recorded robot demonstrations, but it may not be available during deployment.
To overcome the limited data problem, [34, 36, 10] train on a mix of datasets; however, the overall scale remains limited. The total hours reported for ViNT/NoMaD training are $80\text{\,}\mathrm{h}$ across the collected datasets (see Tab. I) and even less for FlowNav, which is modest given the complexity of the task. These works compellingly demonstrate the promise of imitation learning for navigation; nevertheless, the scarcity of expert data continues to hinder generalization.
@@
LiReN [37] approaches data scarcity by continuously improving the model during deployment. They first use offline RL to train a navigation policy on the same multi-embodiment dataset as ViNT. Then they deploy the model and use an actor-critic algorithm to continuously improve it during deployment without needing a human expert to collect more data.
LeLaN [13] uses NoMaD and a vision–language model (VLM) to automatically label first-person view (FPV) YouTube videos for navigation policy learning.
CityWalker [22] also uses a dataset of FPV videos from YouTube, on which it runs visual odometry to automatically generate supervision. The FrodoBots-2k dataset [14] uses crowdsourcing; volunteers remotely drive inexpensive robots (e.g., Earth Rover Zero) to collect demonstrations in urban pedestrian zones. This strategy scales data collection without on-site experts; however, the dataset suffers from poor tracking, inaccurate localization, and low-quality sensors. This is addressed by using Model-Based ReAnnotation (MBRA) [14], an approach that transforms a low-quality dataset into a filtered version with high-quality supervision.
@@
OmniVLA [12] is an omni-modal vision-language-action (VLA) model for navigation that supports goal conditioning via 2D goal poses, egocentric goal images, natural language instructions, and their combinations. Built on top of OpenVLA [19], OmniVLA leverages a large pretrained VLA backbone and is further trained on approximately $9500\text{\,}\mathrm{h}$ of navigation data. This dataset is the combination of [34, 13, 14, 42] datasets from 10 different robotic platforms. Notably, the dataset mixture is heavily skewed towards BDD-V [42], i.e., on-road autonomous driving data. They use a customized MBRA-style [14] approach to generate synthetic, feasible supervision for their target robots on BDD-V.
@@ III-B MPPI Planning
$\min_{\mathbf{a}_{0:N-1}}\quad J(\mathbf{s}_{0},\mathbf{a}_{0:N-1}),$ (1)
where $\mathbf{a}_{0:N-1}$ is a sequence of command velocities $\mathbf{a}_{i}=[v_{i}^{x},v_{i}^{y},\omega_{i}]$, with $v_{i}^{x},v_{i}^{y},\omega_{i}$ representing the linear and angular velocities, respectively, and $J:(\mathbf{s}_{0},\mathbf{a}_{0:N-1})\mapsto c\in\mathbb{R}$ denoting a predefined cost function that encourages goal-reaching and safety, conditioned on the current state $\mathbf{s}_{0}\in\mathrm{SE}(2)$.
MPPI solves the optimization by iteratively sampling a population of command velocity sequences $\mathbf{a}_{0:N-1}$. It then uses $J$ to weight the population members and update the population distribution, preferring low-cost members.
The future states are computed based on the current state $\mathbf{s}_{0}$ and a dynamics model $\mathbf{s}_{i+1}=f(\mathbf{s}_{i},\mathbf{a}_{i})$. The dynamics model $f$ can be learned or classically derived. In our case, we use a simple piecewise-constant velocity model in $\mathrm{SE}(2)$:
@@
$\begin{bmatrix}x_{i+1}\\ y_{i+1}\end{bmatrix}$ $=\begin{bmatrix}x_{i}\\ y_{i}\end{bmatrix}+\mathbf{R}(\theta_{i})\begin{bmatrix}v_{i}^{x}\\ v_{i}^{y}\end{bmatrix}\Delta t,$ (3)
@@
$\begin{split}J(\mathbf{s}_{0},\mathbf{a}_{0:N-1})&=w_{\text{trav}}C_{\text{trav}}(\mathbf{s}_{1:N})+w_{\text{goal}}C_{\text{goal}}(\mathbf{s}_{1:N})\\ &\quad+w_{\text{effort}}C_{\text{effort}}(\mathbf{a}_{0:N-1}).\end{split}$ (4)
@@ III-B1 Traversability Cost ($C_{\text{trav}}$)
We use the convolutional neural network (CNN) from [25, 7] to compute a traversability map $\mathbf{T}$ from the elevation map $\mathbf{E}$. The CNN is trained to provide an embodiment-specific traversability estimate based on the robot’s capabilities. The traversability cost is the sum of the grid cells of $\mathbf{T}$ traversed by the robot along $\mathbf{s}_{1:N}$.
@@ III-B3 Effort cost ($C_{\text{effort}}$)
This is the weighted sum of the linear, lateral, and angular velocities $w_{\text{lin}}|v^{x}_{i}|+w_{\text{lat}}|v^{y}_{i}|+w_{\text{ang}}|\omega_{i}|$ over the command sequence, encouraging efficient motion while discouraging unnecessary sideways motion and excessive turning.
@@ III-C1 Teleoperation Dataset ($\mathcal{D}_{\text{TEL}}$)
$\mathcal{D}_{\text{TEL}}$ consists of fixed-horizon, goal-conditioned segments extracted from the recorded teleoperated trajectories and serves to transfer semantic preferences and motion characteristics to the policy.
@@ III-C2 Geometric Dataset ($\mathcal{D}_{\text{GEO}}$)
To increase the diversity and scale of the training data, we employ the previously introduced MPPI-planner to generate additional demonstrations. For every camera frame $\mathbf{I}$, we independently sample $K$ random goals $\mathbf{g}_{1},...,\mathbf{g}_{K}$ from a Gaussian distribution in the robot-centric coordinate frame (see Tab. III in the Appendix). Given the matching elevation map $\mathbf{E}$ and the goals, MPPI plans the paths $\mathbf{s}_{1:N}^{1},...,\mathbf{s}_{1:N}^{K}$; generating $K$ diverse training samples $(\mathbf{I},\mathbf{g}_{k},\mathbf{s}_{1:N}^{k})$ that are saved in $\mathcal{D}_{\text{GEO}}$.
The MPPI planner reliably generates geometrically feasible paths by leveraging privileged information from accurate elevation mapping. This enables the creation of demonstrations that are more diverse and exploratory than those obtained from real-world teleoperation alone. Even when a randomly sampled goal is infeasible, the planner still produces safe, embodiment-consistent behavior, either progressing toward the goal in a best-effort manner or exploring alternative routes while avoiding obstacles (see Fig. 2).
@@ III-D Policy Architecture
Figure 3: Overview of LiMo’s architecture. The policy takes as input a single RGB image $\mathbf{I}$, and a goal pose $\mathbf{g}=(x,y,\theta)$ in the robot-centric frame. Image features are extracted using a frozen DINOv2 [27] encoder and combined with learned positional embeddings. A transformer decoder conditioned on the goal embedding predicts NN waypoint embeddings, which are linearly projected to $N$ waypoints $(x,y,\theta)$, forming the output trajectory.
@@ IV-A Datasets
The experiments build on the GrandTour dataset [9], which comprises $6\text{\,}\mathrm{h}$ of highly diverse deployments of an ANYmal D robot across 49 sites in Switzerland, with varying weather and lighting conditions. From these missions, we derive three datasets, in accordance with Secs. III-C1 and III-C2: (i) the teleoperation dataset $\mathcal{D}_{\text{TEL}}$, (ii) the geometric-planner-labeled dataset $\mathcal{D}_{\text{GEO}}$ with $K=10$ trajectories per image, each with $N=50$ waypoints evenly spaced over $T=$ $5\text{\,}\mathrm{s}$, and finally, (iii) the augmented dataset $\mathcal{D}_{\text{AUG}}=\mathcal{D}_{\text{TEL}}\cup\mathcal{D}_{\text{GEO}}$. For evaluation, we reserve six GrandTour missions exclusively as test scenes (see Tab. III in the Appendix for further details on dataset size).
@@ IV-C1 Metrics
Following the recommendations of [3], we evaluate trajectories on held-out missions using geodesic distances (GD) and Success weighted by normalized inverse Path Length (SPL). A path is deemed successful if (i) the GD of the final waypoint to the goal is $\leq$1.0\text{\,}\mathrm{m}$\$ and (ii) the robot does not collide. Collisions and geodesic distances are computed using traversability maps derived from the ground-truth elevation maps. Let $S_{i}\in\{0,1\}$ indicate success for sample $i$, $\ell_{i}$ be the geodesic start-to-goal distance, and $p_{i}$ the path length; then the SPL is defined as:
$\mathrm{SPL}=\frac{1}{M}\sum_{i=1}^{M}S_{i}\,\frac{\ell_{i}}{\max(p_{i},\ell_{i})}\,.$ (5)
@@ IV-C2 Baselines
Geometric Planner Oracle
The path obtained by running the planner described in Sec. III-B. The geometric planner should be viewed as an approximate oracle rather than a deployable baseline. It plans with access to the ground-truth elevation maps that are not available to the learned policies.
Moreover, the planner is used to define the evaluation set: we consider only goals that it can reach, so as not to penalize methods for inherently infeasible targets.
Straight-Line Paths
Paths going straight from the robot to the goals, capped at $5\text{\,}\mathrm{m}$ in length.
Real-World Paths
The real-world paths recorded during dataset collection, only available for $\mathcal{D}_{\text{TEL}}$ goals.
@@ IV-C3 Results
OmniVLA (open-loop) $12.6$ $67.2$ $41.5$
@@
Straight-Line Paths $12.5$ $87.5$ $87.5$
Real-World Paths $11.0$ $89.0$ $87.5$
@@
OmniVLA (open-loop) $14.2$ $46.5$ $33.5$
@@
Straight-Line Paths $23.1$ $76.9$ $76.9$
@@
As shown in Tab. II, incorporating geometric demonstrations yields substantial SPL improvements on the diverse $\mathcal{D}_{\text{AUG}}$ test set, indicating that geometric supervision significantly strengthens performance in complex, unseen scenarios (Q1, Q2).
@@
The non-zero collision rate of the Real-World Paths baseline mainly reflects limitations of the traversability map based collision proxy. Collisions are determined from elevation maps, which can label rough but physically traversable terrain as obstacles. As a result, safely executed recordings are sometimes counted as collisions. This makes the reported collision rates conservative and biased toward overestimating failures.
We observe that OmniVLA achieves competitive collision avoidance, even on the more challenging $\mathcal{D}_{\text{AUG}}$ test set, but exhibits reduced goal-reaching performance. This is expected in part due to the evaluation protocol: OmniVLA is trained for closed-loop replanning with fixed-length waypoint chunks rather than open-loop trajectory termination at a specified goal.